# VLM-Assisted Continual learning for Visual Question Answering in Self-Driving

### v1

@@ 4.1 Experimental Settings
Compared Methods. We compare our method with the following methods on the DriveLM dataset [29]: the weight regularization method EWC [14], the function regularization method LwF [17], and the memory replay method DER [4] to validate the effectiveness of our approach. In addition, we provide a lower bound (Vanilla), which simply performs gradient updating without any countermeasures against forgetting, and a relative upper bound (Joint), which trains all tasks jointly. Moreover, we compare another method that uses a joint training approach but is based on a different model, referred to as DriveLM-Agent [29], to demonstrate the performance differences in handling multitask learning across various model architectures.
Implementation Details.Our implementation is based on the EM-VLM4AD framework [11], utilizing a Vision Transformer (ViT) [8] for multi-view image embedding and a pre-trained T5 model for text processing, enabling Visual Question Answering (VQA) in autonomous driving. We enhance this with a projection layer for better embedding integration and employ continual learning strategies like memory replay and knowledge distillation to mitigate catastrophic forgetting and retain knowledge across tasks. We implement our model based on Pytorch [22] on a single NVIDIA L20 GPU. The training phase of each task is performed for 4 epochs, and each model takes about 1 day to complete. For hyper-parameters, we use a learning rate of 1e-4, a weight decay of 0.05, a batch size of 4, a temperature parameter set to 2 and $\alpha$ value is set to 0.7 for the knowledge distillation part. A dynamic weight parameter to adjust the projection regularization loss, where the $\lambda$ parameter is initialized to 0.05 and halved progressively with each task. More details of experimental hyperparameter settings and experiments refer to supplementary materials.
4.2 Results
Quantitative Analysis. As shown in Table 1, our model significantly outperforms existing continual learning methods in key performance indicators for language and vision tasks. Our approach yielded the highest scores in BLEU, METEOR, ROUGE, and CIDEr metrics. Specifically, our method achieved a BLEU-1 score of $69.85\%$, surpassing the next best method (Joint training) by $2.13\%$. In more complex metrics, such as BLEU-4 and METEOR, our model demonstrated improvements of $5.75\%$ and $4.10\%$ points over the Joint method, respectively, indicating a better ability to handle nuanced language generation in the context of autonomous driving and highlight our model’s enhanced ability to generate context-aware descriptions of the driving environment, crucial for interpreting dynamic situations like traffic signals, pedestrians, and obstacles. The improvement in CIDEr scores from 3.2 to 3.4 also suggests enhanced capability in generating relevant, human-aligned descriptions of images, improving the model’s effectiveness in interpreting and reacting to dynamic driving environments. Our experiments demonstrate the efficacy of our proposed continual learning VLM model for autonomous driving across various critical metrics.
Furthermore, as illustrated in Table 1, our model outperforms traditional continual learning methods like EWC, LwF, and DER across several metrics. Specifically, it surpasses EWC by $15.66\%$ and LwF by $21.25\%$ in BLEU-4 scores, highlighting the limitations of methods relying solely on regularization. Unlike these approaches, our model incorporates memory replay and advanced knowledge distillation, improving task retention and prediction refinement. Additionally, while DER includes memory replay, it lacks projection layer regularization and knowledge distillation, key components in our method for handling the complexities of multimodal data in autonomous driving. Our model also improves METEOR and ROUGE_L scores by $1.34\%$ and $3.25\%$ over DER. Moreover, our CIDEr score of 3.4 is significantly better than all methods we compare. This highlights the effectiveness of our dynamic projection regularization, which constrains model over-migration through regularization. These comparative results suggest that persistence is crucial in situations where models encounter diverse and unpredictable environments, ensuring high accuracy and robustness under different driving conditions.
Figure 3: Example of correct and failure answer generations from our model.
@@ 4.3 Ablation Studies
Memory Reply and Knowledge Distillation. Using memory replay/experience replay (ER) alone significantly enhances the model’s continual learning performance. While knowledge distillation (KD) has a limited effect on traditional metrics, it effectively improves the CIDEr score by smoothing probability distributions, thereby reducing overconfident predictions and enhancing prediction stability.
@@
4.4 Analysis
Performance of Our Method on Other VQA Datasets. The table 3 illustrates the effectiveness of our proposed method on the Nuscenes-QA [24] dataset, demonstrating significant improvements across critical evaluation metrics. With the incorporation of continual learning (CL), our approach achieves notably higher accuracy in both the ’Exist’ and ’Comparison’ tasks, accompanied by a substantial increase in the overall ’Average’ score. Additionally, the marked reduction in the ’Forget’ rate underscores the model’s enhanced capability for knowledge retention and adaptability in evolving environments.
Adaptability of CL method with Other VLM model.We replaced the original ViT and T5 backbones in our VLM framework with LLaMA-Adapter-v2[10] to evaluate the adaptability of our continual learning (CL) approach. Table 4 compares models without CL (”w/o CL”) and with CL (”w/ CL”) using LLaMA-Adapter-v2, showing significant improvements in BLEU, METEOR, ROUGE_L, and CIDEr scores. This comparison aims to evaluate the robustness and adaptability of our continual learning (CL) approach when applied to alternative model architectures. From the results, the BLEU, METEOR, ROUGE_L, and CIDEr scores improved significantly, demonstrating enhanced fluency, contextual accuracy, and alignment with human references, particularly for image-related tasks. This comparison further underscores the importance of continual learning in maintaining performance in dynamic and multi-tasks autonomous driving scenarios. It also highlights that our method is not limited to specific backbones, offering a flexible solution for integrating continual learning into different vision-language frameworks.
Method BLUE-1 BLUE-2 BLUE-3 BLUE-4
w/o CL 49.29 44.55 41.36 38.74
w/CL 67.41 62.41 58.08 54.05
Method METEOR ROUGE_L CIDEr
w/o CL 28.72 61.40 2.37
w/CL 38.87 73.28 3.17
@@ 4.5 Case Analysis
Figure 3 illustrates our approach to multitask learning with the autonomous driving framework, dividing the outcomes into correct and failure cases. In correct cases, for example, after the model is trained across multiple tasks, the application of continual learning methods enables it to retain its ability to accurately perform the Perception task. The model correctly answers queries about the status of a motorcycle, identifying it as without a rider. However, when continual learning methods are not employed, the model’s responses are dominated by more recent tasks, causing it to lose crucial information from earlier tasks, highlighting the adverse effects of catastrophic forgetting in multitask scenarios.
Even with continual learning, the model struggles in failure cases, likely due to dataset distribution imbalance, such as in the Behavior task where the training examples of parked cars are very limited. Further case analyses are provided in the supplementary materials..
@@ 5 Conclusion
In this paper, we presented a continual learning approach applied to vision language models (VLMs) for autonomous driving with VQA task. The method combines memory replay and knowledge distillation, and demonstrates its strong performance on the DriveLM dataset by introducing a projection layer for regularization in feature embedding. Extensive experimental results demonstrate the effectiveness and generalizability of the proposed model.

### v2

@@ 4.2 Analysis
4.3 Quantitative Results
As shown in Table 1, our model significantly outperforms existing continual learning methods across all key metrics, achieving the highest BLEU, METEOR, ROUGE, and CIDEr scores. Specifically, it reaches a BLEU-1 score of $71.25\%$, surpassing the next best method (Joint) by $4.43\%$, and improves BLEU-4 and METEOR by $8.04\%$ and $5.17\%$, respectively, demonstrating stronger capability for nuanced, context-aware language generation in autonomous driving. The CIDEr score improvement from 3.2 to 3.4 further confirms the model’s ability to produce relevant, human-aligned image descriptions. Our model also outperforms EWC, LwF, and DER, exceeding them by $17.95\%$, $23.54\%$, and $4.33\%$ in BLEU-4, respectively. These results highlight the limitations of regularization approaches and the benefits of combining memory replay with advanced knowledge distillation and projection-layer regularization, which enable better task retention and multimodal understanding. Moreover, our CIDEr score of 3.4, the highest among all methods, demonstrates the effectiveness of dynamic projection regularization in preventing feature drift and maintaining robustness across driving scenarios.
@@ 4.4 Ablation Studies
Memory replay and Knowledge Distillation. Using memory replay/experience replay (ER) alone significantly enhances the model’s continual learning performance. While knowledge distillation (KD) has a limited effect on traditional metrics, it effectively improves the CIDEr score by smoothing probability distributions, thereby reducing overconfident predictions and enhancing prediction stability.
@@
Method BLEU-4 METEOR ROUGE_L CIDEr
w/o CL 38.74 28.72 61.40 2.37
w/CL 54.05 38.87 73.28 3.17
@@
Figure 3: Examples of correct and failure answer generations from our model.
Performance of Our Method on Other VQA Datasets. The table 3 illustrates the effectiveness of our proposed method on the Nuscenes-QA [nuscenesqa] dataset, demonstrating significant improvements across critical evaluation metrics. With the incorporation of continual learning (CL), our approach achieves notably higher accuracy in both the ’Exist’ and ’Comparison’ tasks, accompanied by a substantial increase in the overall ’Average’ score. Additionally, the marked reduction in the ’Forget’ rate underscores the model’s enhanced capability for knowledge retention and adaptability in evolving environments.
Adaptability of CL method with Other VLM model.
To evaluate the adaptability of our continual learning (CL) approach, we replaced the backbone in our VLM framework with LLaMA-Adapter-v2 [llama]. As shown in Table 4, applying CL (”w/ CL”) improves BLEU, METEOR, ROUGE_L, and CIDEr scores over the non-CL baseline (”w/o CL”). These gains reflect enhanced fluency, contextual accuracy, and alignment with human references in image-grounded tasks. This result demonstrates the robustness of our CL method across architectures and highlights its flexibility for integration into diverse vision-language frameworks, especially in autonomous driving scenarios.
@@ 4.5 Case Analysis
Figure 3 presents examples of correct and failure cases in our multitask autonomous driving framework. Note that each question in the figure is evaluated at four different stages of continual learning. After completing task $T_{n}$, the model is re-evaluated on all previously learned tasks; therefore, a single perception question yields four answers corresponding to the checkpoints obtained after training Perception, Prediction, Planning, and Behavior. This setup allows us to visualize whether the model preserves earlier knowledge as new tasks are introduced.
In the correct examples, with joint training on perception, prediction, planning, and behavior tasks—and the use of continual learning methods—the model maintains its accuracy on the perception task. For instance, in the pedestrian scenario, it consistently answers “Two pedestrians are moving,” and in the trailer scenario, it consistently responds “The trailer in front of the ego car is parked,” across perception, prediction, planning, and behavior stages. By contrast, without continual learning, the model’s outputs become dominated by information from later tasks, causing crucial details from earlier perception subtasks to be lost and illustrating catastrophic forgetting in a multitask setting.
Despite continual learning, data imbalances still cause errors. For example, in the behavior task, scarce parked‑vehicle samples lead the model to predict “Going ahead” when the vehicle has braked, or to mislabel a black sedan as “Changing to the left lane” or “No entry” in trailer scenarios. These failure cases underscore the challenge of rare samples—see the supplementary materials for further analysis.
@@ 5 Conclusion
In this paper, we presented a continual learning approach applied to vision language models (VLMs) for autonomous driving with VQA task. The method combines memory replay and knowledge distillation, and demonstrates its strong performance on the DriveLM dataset by introducing a projection layer for regularization in feature embedding. Extensive experimental results demonstrate the effectiveness and generalizability of the proposed model. In future work, we will explore more dynamic continual learning strategies to further improve model robustness.
[…]